Shift Summary Evening 9 July 1999

Douglas A. Jensen (djensen@fnal.gov)
Fri, 9 Jul 1999 23:58:03 -0500 (CDT)

Shift Summary for Evening Shift, 09/Jul/99

On Shift: Viv O'Dell, Doug Jensen
Also Miguel, Marj, Sasha 3, Ed B.(in person, by phone), Rick K (phone)

Hours of beam available at all intensities: 6.4
Hours of beam available at nominal intensity: 6.4, but nom. redefined
Hours of E832 data (special runs not included): 0.
Hours of special runs (laser scans, muon runs...): 0.

Accounting of lost time:

F1 Quench @ 4:23 to A3 quench to recovery at about 6:25. (2hrs)
DAQ Computer system problems during this time. No extra lost time.

A reset by power cycle, of laser crate.

Tape change: none

========================================================================
Runs ended this shift:
Run 13844 - F1 quench, atev? mucking, sdaq repairs ?

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:

Paul Czarapata and John Mariner stopped by to chat.

Running at beginning, then F1 quench while sdaq problems, Viv discussed
with Theo,.... Turns out it was a security upgrade that wiped out
the system. After some phone calls, Viv fixed it all.

After recaovery, changed to High Intensity running, first requesting
7.0 E 12, but then changed the request to 6.8E12 in order to not run
out of CPU resources ( clean up the live time, see log ).

11:45 hand over to night shift - who have so kindly arrived early.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13844 16:00 new shift crew
16:40 run has definately ended.
.... 17:55 Awaiting recovery from A3 quench, expet beam
between 18:00 and 18:15
13845 1 18:25 Beam reappears, starting at about 1.5E12, soon
up to 3.7E12. BOA at about 1.3.
20 18:50 Intensity up. 7.33E12, Boa 1.09.
93 20:27 Viv runs KFRIEND, Rick K decided that the amount of
data going to plane3 needed reduction. tweeked.
Marj notes that DC3y draws about 1.8 microamps off
spill, other planes much less.
106 20:45 Lower request, as per EB, to 6.8E12, to stay below
7\E12. Above 7E12 we get hammered with dead time
due to the spill not fully analyzed before the next
spill starts (see logbook). Lowering intensity
request fixes the problem.
191 10:40 KFRIEND - stream 5 problem. Worked
241 11:45 KFRIEND
Night shift guys recycle power to laser crate.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift
No muon run. Keep taking high intensity data.

Douglas Jensen DJensen@fnal.gov x8194